OpenStudy (anonymous):

Water freezes at 0 degrees. If water at 97 degrees is cooled at a rate of 6 degrees per min, will it be forzen in 16 min? show a mathematical statement and all of your work . . . help! @OpenStudy Feedbac…

OpenStudy (saifoo.khan):

it will freeze. :P

OpenStudy (anonymous):

ok, and why>

OpenStudy (anonymous):

i need to say why lol

OpenStudy (anonymous):

97-6(x)=0, Where x is the number of minutes for the water to reach 0 degrees -6(x)=-97 6x=97 x=97/6 x=16 and 1/6minutes

OpenStudy (shadowfiend):

Please don't tag OpenStudy Feedback on questions that are not related to the OpenStudy site. Thanks!